# Query-Based Adaptive Aggregation for Multi-Dataset Joint Training Toward Universal Visual Place Recognition

### v1

@@ Introduction
Datasets. For multi-dataset joint training, we utilize GSV-Cities (Ali-bey, Chaib-draa, and Giguère 2022) with sparse place sampling, MSLS (Warburg et al. 2020) incorporating place cluster information, and SF-XL (Berton, Masone, and Caputo 2022), which is grouped by location and orientation. These large-scale datasets encompass a wide range of variations essential for outdoor VPR, as outlined in Table 1.
@@
For evaluation, we incorporate a wide range of VPR datasets. The AmsterTime dataset (Yildiz et al. 2022) matches historical grayscale images with current RGB images, testing robustness to long-term temporal changes. The Eynsham dataset (Cummins and Newman 2009) focuses on grayscale images for VPR. Pittsburgh (Torii et al. 2013), Tokyo24/7 (Torii et al. 2015), and SF-XL (Berton, Masone, and Caputo 2022) datasets utilize Google Street View as the source of database images and primarily assess viewpoint variations. MSLS (Warburg et al. 2020) collects images from a forward-facing car-mounted camera. Nordland∗∗(Sünderhauf, Neubert, and Protzel 2013) examines seasonal changes between summer and winter conditions, using a 10-frame threshold for positive samples, while its variant, Nordland(Zaffar et al. 2021), employs a 1-frame threshold, demanding higher accuracy. The SPED dataset (Chen et al. 2017) captures seasonal and day-night variations using surveillance footage. SVOX (Berton et al. 2021) poses varying illumination and weather conditions.
@@
Evaluation Metrics. We evaluate VPR performance using Recall@1 (R@1), focusing on the proportion of query samples with correct top-1 matches. Following prior works (Ali-bey, Chaib-draa, and Giguère 2024; Izquierdo and Civera 2024b), positive matches are defined using a 25 m distance threshold for most datasets, while frame-based thresholds are used for datasets like Nordland (Sünderhauf, Neubert, and Protzel 2013; Zaffar et al. 2021).
@@
Implementation details. For multi-dataset joint training, we set $k=30$ and $m=4$ per dataset, resulting in a batch size of $90$ places and $360$ images. Training images are resized to $224\times 224$ and evaluation images to $322\times 322$. We fine-tune the last two blocks of the DINOv2 model (Oquab et al. 2024) alongside the QAA module, selecting models based on MSLS val performance. We optimize models using the AdamW optimizer (Loshchilov and Hutter 2019) with a learning rate of $4\times 10^{-5}$ for aggregation layers. Training spans at most 80 epochs, with approximately 2000 iterations per epoch and a 4000-iteration linear learning rate warmup. A weight decay of $1\times 10^{-3}$ is applied. Clique Mining (CM) (Izquierdo and Civera 2024a) is used to group place data, with cliques being recomputed during training. The implementation is based on PyTorch Lightning, and the training is conducted on a single NVIDIA A100 GPU.
Results
Comparison with Baselines
@@ Results
In this section, we evaluate the R@1 performance of our proposed QAA method in comparison with state-of-the-art VPR methods. Results for multi-view datasets are presented in Table 2, while those for front-view datasets are shown in Table 3. Our comparison includes leading baselines such as NetVLAD (Arandjelovic et al. 2016), SRFS (Ge et al. 2020), Conv-AP (Ali-bey, Chaib-draa, and Giguère 2022), MixVPR (Ali-bey, Chaib-draa, and Giguère 2023), CosPlace (Berton, Masone, and Caputo 2022), EigenPlace (Berton et al. 2023), BoQ (Ali-bey, Chaib-draa, and Giguère 2024), and SALAD with Clique Mining (CM) (Izquierdo and Civera 2024b, a). These methods represent some of the most effective and widely recognized one-stage approaches in the field. Notably, recent baselines like BoQ and SALAD CM have demonstrated superior accuracy and efficiency than two-stage methods.
@@
Table 4: Performance Comparison for Joint Training across Different Datasets with SALAD CM (Izquierdo and Civera 2024a) and Our QAA Approaches ($N_{q}=256,C_{f}=64,C_{r}=128,C_{d}=8192$). All methods incorporate Clique Mining (CM) (Izquierdo and Civera 2024a) and DINOv2-B backbone. The best results are highlighted in bold.
@@
Tables 2 and 3 reveal distinct strengths for BoQ and SALAD CM on multi-view and front-view datasets, respectively, due to their specialized training. BoQ’s training on a multi-view dataset creates a bias for such data, while SALAD CM’s training on both, combined with a limited model capacity, leads to overfitting on front-view characteristics. Our QAA method addresses these limitations by efficiently enhancing information capacity and extensive joint training. This strategy enables QAA to learn more generalized representations that are effective across both multi-view and front-view datasets, demonstrating superior generalization capabilities compared to BoQ and SALAD CM.
@@
Specifically, as presented in Table 2, the QAA method surpasses BoQ on the AmsterTime, Eynsham, Pitts30k, SF-XL v1, and Tokyo 24/7 datasets, which are multi-view in nature. On the Pitts250k, SPED, and SF-XL v2 datasets, QAA achieves performance comparable to BoQ despite utilizing a significantly smaller output dimension (8192 vs. 12288). The reduced output dimension not only indicates computational efficiency but also suggests that QAA captures essential features without the need for a larger output descriptor dimension. These results highlight the effectiveness of our method on multi-view datasets and underscore its competitive advantage over state-of-the-art baselines.
@@
Similarly, as shown in Table 3, QAA significantly outperforms both SALAD CM and BoQ on the MSLS and Nordland evaluation sets. It also achieves performance on par with SALAD CM and BoQ on the SVOX evaluation sets. Notably, despite SALAD CM’s training bias toward front-view datasets, our method delivers superior overall performance on front-view datasets. This demonstrates that QAA yields balanced and optimal performance to different dataset types, validating its robustness and adaptability.
@@
Performance of Reduced $C_{d}$. Tables 2 and 3 present the performance of reduced $C_{d}$. For $C_{d}=4096$ and $2048$, QAA overall maintains strong performance, particularly on the Eynsham, Pitts250k, Pitts30k, SPED, SF-XL, Tokyo24/7, MSLS Val and Challenge, Nordland, and SVOX evaluation sets. A mild performance degradation is observed for $C_{d}=1024$ on AmsterTime, SF-XL v1, MSLS Challenge, Nordland, SVOX Night, and SVOX Sun, while performance remains competitive on other datasets. These results highlight the robustness of QAA.
@@
Inference Complexity. We compare the computational and parameter complexity of DINOv2 SALAD, BoQ, and QAA. SALAD employs 1.4M parameters for its aggregation module and, for a $322\times 322$ image, requires 0.94 GFLOPS with convolution-based aggregation. For attention-based aggregation, BoQ, with 64 queries, utilizes 8.6M parameters and 8.22 GFLOPS. Compared with BoQ, QAA, despite using 256 queries, requires only 5.1M parameters and 2.29 GFLOPS while maintaining the reduced output size.
Table 5: Performance Comparison for Joint Training across Different Settings for Aggregation Methods with Independent Reference Codebooks. The best results are in bold.
Methods MSLS Val Pitts250k Val SF-XL Val Nordland AmsterTime
Softmax 97.0 94.9 97.7 90.5 61.5
OT 97.3 95.3 97.5 90.9 62.0
CS 97.6 95.4 97.8 91.8 63.7
@@
Table 6: Performance Comparison for Joint Training across Different Numbers of Queries $N_{q}$ with $C_{f}=64$ and $C_{r}=128$. The best results are highlighted in bold. For inference complexity, GFLOPS is computed for the QAA module using a single $322\times 322$ image.
@@
Ablation Study
We evaluate the effectiveness and examine the underlying mechanism of our QAA approach, aiming to achieve balanced and optimal performance across a wide range of datasets. We assess validation R@1 performance on MSLS, Pitts250, and SF-XL, as their performance is closely tied to the training datasets. To evaluate generalization performance across markedly different data attributes, we include the Nordland and AmsterTime datasets in our study. These datasets introduce significant domain variations, such as seasonal transitions from summer to winter and the presence of historical grayscale images.
Cross-Dataset Evaluation.
Table 4 presents the cross-dataset evaluation results using the DINOv2 SALAD CM approach (Izquierdo and Civera 2024a) and our QAA approaches. The findings reveal that models trained on individual datasets achieve the highest accuracy on their respective evaluation sets but struggle on others, indicating an inductive bias toward the specific characteristics of the training data. In contrast, joint training on GSV-Cities, MSLS, and SF-XL leads to more balanced performance across all three validation sets, as well as on significantly different datasets such as Nordland and AmsterTime. However, joint training negatively impacted MSLS Val performance on SALAD CM, though QAA still performed comparably to models trained exclusively on MSLS. Moreover, our QAA approaches consistently outperform the SALAD CM method across various training configurations, with particularly strong improvements on MSLS Val, SF-XL Val, Nordland, and AmsterTime, underscoring QAA’s superior accuracy.
@@
Effectiveness of Independent Reference Codebooks and CS Matrix.
Table 5 presents a performance comparison of QAA under different aggregation paradigms: (1) Softmax, (2) Optimal Transportation (OT), and (3) Cross-query Similarity (CS). The results highlight two key observations: (1) the efficacy of using an independent reference codebook across different paradigms, and (2) the superior performance of CS, particularly on the MSLS Val, Nordland, and AmsterTime datasets, emphasizing the advantage of similarity-based aggregation in enhancing VPR representations.
Coding Rate Analysis.
We analyze the information capacity of $\hat{P}$ by examining the histogram of coding rates per query using the MSLS Val in Fig. 4. The results indicate that $\hat{P}$ generated by CS exhibits a 2$\sim$ 2$\times$ 2 × coding rate with reduced variance compared to Softmax and OT, both of which compress the output space and restrict information content. In contrast, CS preserves more information in query-level image features, enabling richer interactions with the reference codebook. This enhanced information retention facilitates the generation of highly informative descriptors, ultimately leading to superior performance.
Query Scalability.
Table 6 explores the relationship between the number of queries, $N_{q}$, and the performance of QAA. The results indicate that increasing $N_{q}$ enhances QAA’s performance on Pitts250k, SF-XL Val, Nordland, and AmsterTime, with performance gains saturating at $N_{q}=128$ and $N_{q}=256$. The MSLS Val performance remains stable across different $N_{q}$, indicating that this evaluation set follows a consistent pattern from the fixed front-viewpoint. Therefore, fewer queries are required to capture the pattern effectively. This highlights the importance of a larger $N_{q}$ in optimizing QAA’s effectiveness.
Effect of Channel Numbers.
Table 7 analyzes performance variations across different values of $C_{f}$ and $C_{r}$, which influence the final output dimension $C_{d}$. We observe that reducing either $C_{f}$ or $C_{r}$ results in a slight but comparable performance degradation. Interestingly, the model remains robust even when $C_{f}$ is extremely reduced to $8$, despite the information bottleneck in the query-level image feature $\hat{P}$, whose dimension shrinks to $256\times 8$. This resilience is attributed to the support of the high-dimensional codebook $\hat{F}$, which helps maintain performance stability.
Table 7: Performance Comparison for Joint Training across Different Numbers of Channels for Learned Queries $C_{f}$ and $C_{r}$ with $N_{q}=256$. The best results are highlighted in bold.
@@
  • Global Context Capture: The multi-attention module allows attention maps to span the entire image, capturing global context rather than relying solely on patch-level feature combinations. This broader perspective likely contributes to the performance gains in experiments.
  • Diverse Attention Patterns: Different query vectors exhibit unique attention patterns. For instance, some queries focus on distant objects in the foreground, while others emphasize nearby roads or structures. This diversity in attention may further enhance model performance.

### v4

@@ IV Experimental Setup
Datasets. For multi-dataset joint training, we utilize GSV-Cities [1] with sparse place sampling, MSLS [39] incorporating place cluster information, and SF-XL [5], which is grouped by location and orientation. These large-scale datasets encompass a wide range of variations essential for outdoor VPR, as outlined in Table I.
@@
For evaluation, we incorporate a wide range of VPR datasets. The AmsterTime dataset [42] matches historical grayscale images with current RGB images, testing robustness to long-term temporal changes. The Eynsham dataset [13] focuses on grayscale images for VPR. Pittsburgh [33], Tokyo24/7 [32], and SF-XL [5] datasets utilize Google Street View as the source of database images and primarily assess viewpoint variations. MSLS [39] collects images from a forward-facing car-mounted camera. Nordland∗∗[31] examines seasonal changes between summer and winter conditions, using a 10-frame threshold for positive samples, while its variant, Nordland[44], employs a 1-frame threshold, demanding higher accuracy. The SPED dataset [11] captures seasonal and day-night variations using surveillance footage. SVOX [8] poses varying illumination and weather conditions.
@@
Evaluation Metrics. We evaluate VPR performance using Recall@1 (R@1), focusing on the proportion of query samples with correct top-1 matches. Following prior works [3, 18], positive matches are defined using a 25 m distance threshold for most datasets, while frame-based thresholds are used for datasets like Nordland [31, 44].
@@
Implementation details. For multi-dataset joint training, we set the number of places per batch to $k=30$ and the number of images per place to $m=4$ for each dataset, resulting in a batch size of $90$ places and $360$ images. Training images are resized to $224\times 224$ and evaluation images to $322\times 322$. We fine-tune the last two blocks of the DINOv2 model [27] alongside the QAA module, selecting models based on MSLS val performance. We optimize models using the AdamW optimizer [23] with a learning rate of $4\times 10^{-5}$ for aggregation layers. Training spans at most 80 epochs, with approximately 2000 iterations per epoch and a 4000-iteration linear learning rate warmup. A weight decay of $1\times 10^{-3}$ is applied. Clique Mining (CM) [17] is used to group place data, with cliques being recomputed during training. The implementation is based on PyTorch Lightning, and the training is conducted on a single NVIDIA A100 GPU, taking about 35 hours.
V Results
V-A Comparison with Baselines
@@
In this section, we evaluate the R@1 performance of our proposed QAA method in comparison with state-of-the-art VPR methods. Results for multi-view datasets are presented in Table II, while those for front-view datasets are shown in Table III. Our comparison includes leading baselines such as NetVLAD [4], SRFS [16], Conv-AP [1], MixVPR [2], CosPlace [5], EigenPlace [9], BoQ [3], and SALAD with Clique Mining (CM) [18, 17]. These methods represent some of the most effective and widely recognized one-stage approaches in the field. Notably, recent baselines like BoQ and SALAD CM have demonstrated superior accuracy and efficiency than two-stage methods.
@@
TABLE IV: Performance Comparison for Joint Training across Different Datasets with SALAD CM [17] and Our QAA Approaches ($N_{q}=256,C_{f}=64,C_{r}=128,C_{d}=8192$). All methods incorporate Clique Mining (CM) [17] and DINOv2-B backbone. The best results are highlighted in bold.
@@
Tables II and III reveal distinct strengths for BoQ and SALAD CM on multi-view and front-view datasets, respectively, due to their specialized training. BoQ’s training on a multi-view dataset creates a bias for such data, while SALAD CM’s training on both, combined with a limited model capacity, leads to overfitting on front-view characteristics. Our QAA method addresses these limitations by efficiently enhancing information capacity and extensive joint training. This strategy enables QAA to learn more generalized representations that are effective across both multi-view and front-view datasets, demonstrating superior generalization capabilities compared to BoQ and SALAD CM.
@@
Specifically, as presented in Table II, the QAA method surpasses BoQ on the AmsterTime, Eynsham, Pitts30k, SF-XL v1, and Tokyo 24/7 datasets, which are multi-view in nature. On the Pitts250k, SPED, and SF-XL v2 datasets, QAA achieves performance comparable to BoQ despite utilizing a significantly smaller output dimension (8192 vs. 12288). The reduced output dimension not only indicates computational efficiency but also suggests that QAA captures essential features without the need for a larger output descriptor dimension. These results highlight the effectiveness of our method on multi-view datasets and underscore its competitive advantage over state-of-the-art baselines.
@@
Similarly, as shown in Table III, QAA significantly outperforms both SALAD CM and BoQ on the MSLS and Nordland evaluation sets. It also achieves performance on par with SALAD CM and BoQ on the SVOX evaluation sets. Notably, despite SALAD CM’s training bias toward front-view datasets, our method delivers superior overall performance on front-view datasets. This demonstrates that QAA yields balanced and optimal performance to different dataset types, validating its robustness and adaptability.
@@
Performance of Reduced $C_{d}$. Tables II and III present the performance of reduced $C_{d}$. For $C_{d}=4096$ and $2048$, QAA overall maintains strong performance, particularly on the Eynsham, Pitts250k, Pitts30k, SPED, SF-XL, Tokyo24/7, MSLS Val and Challenge, Nordland, and SVOX evaluation sets. A mild performance degradation is observed for $C_{d}=1024$ on AmsterTime, SF-XL v1, MSLS Challenge, Nordland, SVOX Night, and SVOX Sun, while performance remains competitive on other datasets. These results highlight the robustness of the resource-efficient QAA variant.
@@
Inference Complexity. We compare the computational and parameter complexity of DINOv2 SALAD, BoQ, and QAA. SALAD employs 1.4M parameters for its aggregation module and, for a $322\times 322$ image, requires 0.94 GFLOPS with convolution-based aggregation. For attention-based aggregation, BoQ, with 64 queries, utilizes 8.6M parameters and 8.22 GFLOPS. Compared with BoQ, QAA, despite using 256 queries, requires only 5.1M parameters and 2.29 GFLOPS, demonstrating its superior efficiency while maintaining a reduced output size.
@@
TABLE V: Performance Comparison for Joint Training across Different Settings for Reference Codebooks and Aggregation Methods. The best results are highlighted in bold. For inference complexity, GFLOPS is computed for the QAA module using a single $322\times 322$ image.
Methods GFLOPS MSLS Val Pitts250k Val SF-XL Val Nordland AmsterTime
Softmax 2.29 97.0 94.9 97.7 90.5 61.5
Softmax - Cond 7.43 97.0 94.9 95.7 89.4 58.0
OT 2.29 97.3 95.3 97.5 90.9 62.0
OT - Cond 7.43 97.0 94.9 96.0 88.1 60.0
CS 2.29 97.6 95.4 97.8 91.8 63.7
TABLE VI: Performance Comparison for Joint Training across Different Numbers of Queries $N_{q}$ with $C_{f}=64$ and $C_{r}=128$. The best results are highlighted in bold. For inference complexity, GFLOPS is computed for the QAA module using a single $322\times 322$ image.
@@
V-B Ablation Study
We evaluate the effectiveness and examine the underlying mechanism of our QAA approach, aiming to achieve balanced and optimal performance across a wide range of datasets. We assess validation R@1 performance on MSLS, Pitts250, and SF-XL, as their performance is closely tied to the training datasets. To evaluate generalization performance across markedly different data attributes, we include the Nordland and AmsterTime datasets in our study. These datasets introduce significant domain variations, such as seasonal transitions and historical grayscale images.
V-B1 Cross-Dataset Evaluation
Table IV presents the cross-dataset evaluation results using the DINOv2 SALAD CM approach [17] and our QAA approaches. The findings reveal that models trained on individual datasets achieve the highest accuracy on their respective evaluation sets but struggle on others, indicating an inductive bias toward the specific characteristics of the training data. In contrast, joint training on GSV-Cities, MSLS, and SF-XL leads to more balanced performance across all three validation sets, as well as on significantly different datasets such as Nordland and AmsterTime. While joint training degrades MSLS Val performance for both models, SALAD CM exhibits a sharper decline, whereas QAA maintains performance comparable to MSLS-only training. Moreover, our QAA approaches consistently outperform the SALAD CM method across various training configurations, with particularly strong improvements on MSLS Val, SF-XL Val, Nordland, and AmsterTime, underscoring QAA’s superior accuracy.
@@ V-B1 Cross-Dataset Evaluation
V-B2 Effectiveness of Independent Reference Codebooks and CS Matrix
Table V compares the performance of QAA across three aggregation paradigms: (1) Softmax, (2) Optimal Transportation (OT), and (3) Cross-query Similarity (CS). The results yield two main insights: (1) employing an independent reference codebook consistently benefits all paradigms, and (2) CS achieves superior performance, particularly on the MSLS Val, Nordland, and AmsterTime datasets. These findings underscore the advantage of similarity-based aggregation in strengthening VPR representations.
We further investigate the use of a conditional reference codebook (Cond) in the Softmax and OT paradigms, aiming to capture inter-image relationships within the codebook. It is worth noting that CS does not support this extension, as incorporating a conditional codebook would reduce it to dual mappings of the same features. Our analysis highlights two key findings: (1) the conditional codebook incurs higher computational cost, since it requires image features as input, and (2) in query-driven mechanisms, it not only fails to enhance performance but instead leads to degradation compared to independent reference codebooks.
V-B3 Coding Rate Analysis
We analyze the information capacity of $\hat{P}$ by examining the histogram of coding rates per query using the MSLS Val in Fig. 4. The results indicate that $\hat{P}$ generated by CS exhibits a $\sim 2\times$ coding rate with reduced variance compared to Softmax and OT, both of which compress the output space and restrict information content. In contrast, CS preserves more information in query-level image features, enabling richer interactions with the reference codebook. This enhanced information retention facilitates the generation of highly informative descriptors, ultimately leading to superior performance.
V-B4 Query Scalability
Table VI explores the relationship between the number of queries, $N_{q}$, and the performance of QAA. The results indicate that increasing $N_{q}$ enhances QAA’s performance on Pitts250k, SF-XL Val, Nordland, and AmsterTime, with performance gains saturating at $N_{q}=128$ and $N_{q}=256$. The MSLS Val performance remains stable across different $N_{q}$, indicating that this evaluation set follows a consistent pattern from the fixed front-viewpoint. Therefore, fewer queries are required to capture the pattern effectively. This highlights the importance of a larger $N_{q}$ in optimizing QAA’s effectiveness.
V-B5 Effect of Channel Numbers
Table VII analyzes performance variations across different values of $C_{f}$ and $C_{r}$, which influence the final output dimension $C_{d}$. We observe that reducing either $C_{f}$ or $C_{r}$ results in a slight but comparable performance degradation. Interestingly, the model remains robust even when $C_{f}$ is extremely reduced to $8$, despite the information bottleneck in the query-level image feature $\hat{P}$, whose dimension shrinks to $256\times 8$. This resilience is attributed to the support of the high-dimensional codebook $\hat{F}$, which helps maintain performance stability.
TABLE VII: Performance Comparison for Joint Training across Different Numbers of Channels for Learned Queries $C_{f}$ and $C_{r}$ with $N_{q}=256$.
@@ V-C Qualitative Results
  • Global Context Capture: The multi-attention module allows attention maps to span the entire image, capturing global context rather than relying solely on patch-level feature combinations. This broader perspective likely contributes to performance gains in experiments.
  • Diverse Attention Patterns: Different query vectors exhibit unique attention patterns. For instance, some queries focus on distant objects in the foreground, while others emphasize nearby roads or structures. This diversity may further enhance model performance.